5.1 Vehicles pass through a junction on a busy road at an average rate of 60 per hour.

5.1.1 Find the probability that none passes in a given 30-minute interval. (4)

5.1.2 What is the expected number passing in two minutes? (2)

5.1.3 Find the probability that this expected number (computed in 5.1.2) actually pass through in

a given five-minute period. (4)

5.2 Experience has shown that 80/200 of all CDs produced by a certain machine are defective. If a

 quality control technician randomly tests twenty CDs, compute each of the following

probabilities:

5.2.1 P (exactly one is defective).

5.2.2 P (at least one CD is defective).

5.2.3 P (no more than two are defective).

5.2.4 Find the mean, variance and standard deviation of the distribution.

(3) (4) (5)



1
Expert's answer
2021-06-30T13:02:34-0400

5.1


"\lambda=60(0.5)=30"

5.1.1

"P(X=0)=\dfrac{e^{-30}\cdot30^0}{0!}=e^{-30}\approx10^{-13}\approx 0"

5.1.2


"E(X)=60(\dfrac{2}{60})=2"

5.1.3


"\lambda=60(\dfrac{5}{60})=5"

"P(X=2)=\dfrac{e^{-5}\cdot5^2}{2!}=12.5e^{-5}\approx0.084224"

5.2

"p=0.4, q=1-p=1-0.4=0.6, n=12"

5.2.1


"P(X=1)=\dbinom{12}{1}(0.4)^1(0.6)^{12-1}"

"=0.01741425869"

5.2.2


"P(X\geq1)=1-P(X=0)"

"=1-\dbinom{12}{0}(0.4)^0(0.6)^{12-0}="

"=0.99782321767"

5.2.3


"P(X\leq2)=P(X=0)+P(X=1)+P(X=2)"

"=\dbinom{12}{0}(0.4)^0(0.6)^{12-0}+\dbinom{12}{1}(0.4)^1(0.6)^{12-1}"

"+\dbinom{12}{2}(0.4)^2(0.6)^{12-2}="

"=0.08344332288"

5.2.4


"\mu=np=12(0.4)=4.8"

"\sigma^2=npq=12(0.4)(0.6)=2.88"

"\sigma=\sqrt{\sigma^2}=\sqrt{2.88}\approx1.697"